You like the movie “Tumbbad” and want to visit the movie as many times as possible. The total satisfaction derived from watching movie is given in the following table. No of times watch the Movie 1st 2nd 3rd 4th 5th 6th Total Satisfaction 400 700 950 1170 1270 1300 If the cost of watching movie each time is Rs. 220/-, then how many times you will watch theMovie?


1
Expert's answer
2021-02-11T04:52:40-0500

Let's calculate the marginal utility and marginal utility per Rs using the formulas below:


"MU=\dfrac{TU_f-TU_i}{Q_f-Q_i},""MU_{per\ Rs}=\dfrac{MU}{P}."


After the calculations we get the next table:



As we can see from the table, the maximum marginal utility per Rs corresponds to the first time movie watched. Therefore, you will watch the movie one time only.